### v1

# DiCoFlex: Model-agnostic diverse counterfactuals with flexible control

Abstract
Counterfactual explanations play a pivotal role in explainable artificial intelligence (XAI) by offering intuitive, human-understandable alternatives that elucidate machine learning model decisions. Despite their significance, existing methods for generating counterfactuals often require constant access to the predictive model, involve computationally intensive optimization for each instance and lack the flexibility to adapt to new user-defined constraints without retraining. In this paper, we propose DiCoFlex, a novel model-agnostic, conditional generative framework that produces multiple diverse counterfactuals in a single forward pass. Leveraging conditional normalizing flows trained solely on labeled data, DiCoFlex addresses key limitations by enabling real-time user-driven customization of constraints such as sparsity and actionability at inference time. Extensive experiments on standard benchmark datasets show that DiCoFlex outperforms existing methods in terms of validity, diversity, proximity, and constraint adherence, making it a practical and scalable solution for counterfactual generation in sensitive decision-making domains.
1 Introduction
Counterfactual explanations (CFs) have become an integral part of explainable artificial intelligence (XAI) by providing human-interpretable insights into complex machine learning models [10]. CFs answer critical "what-if" questions by suggesting minimal and meaningful changes to input data that could alter the outcome of a predictive model [40]. Such explanations have valuable applications across sensitive domains, including finance, healthcare, and legal decisions, where understanding the predictions of the model and potential alternative outcomes is paramount [42, 9].
(a) Unconstrained
(b) Sparsity constraints
(c) Actionability constraints
Figure 1: Visualization of diverse counterfactual explanations generated by DiCoFlex for a single instance using an artificial two-class moon-shaped dataset. Each subfigure demonstrates different constraint scenarios: (a) no constraints applied, showing natural diversity; (b) sparsity constraints enforced through the $p$-norm parameter, resulting in minimal feature modifications; and (c) actionability constraints applied via feature masks, restricting which features can be modified.
According to Guidotti [10], an ideal counterfactual explanation should satisfy several key properties. First, it must demonstrate validity by successfully changing the model’s prediction. Second, it should maintain proximity by remaining as close as possible to the original input, minimizing the amount of change. Third, it should exhibit sparsity by modifying only the smallest possible set of features. Fourth, it must ensure plausibility by remaining within the realm of realistic, observed data. Finally, it should guarantee actionability by suggesting changes that are feasible for the stakeholder to implement.
In practice, however, these criteria often conflict. For instance, the smallest possible change may result in samples outside the realistic data distribution. To address this, Mothilal et al. [19] and Laugel et al. [17] argue that, rather than producing a single “best” counterfactual explanation, a diverse set of CFs should be generated. Presenting multiple realistic alternatives allows users to see different ways to achieve the desired outcome, explore trade-offs between minimal change and realism, and gain deeper insight into the model’s decision boundaries. This diversity, in turn, empowers stakeholders to make more informed, feasible decisions.
Current approaches to diverse counterfactual generation face significant limitations. Optimization-based methods like DiCE [19] and multi-objective frameworks [5] incorporate diversity directly into their cost functions but require solving complex, separate optimizations for each explanation. This results in high computational costs and latency, making them impractical for real-time applications. Meanwhile, generative approaches such as those by Pawelczyk et al. [25] and Panagiotou et al. [22] offer efficiency improvements but frequently produce redundant explanations, require access to model gradients, and lack flexibility for adapting to user-defined constraints without extensive retraining.
To address these limitations, we introduce DiCoFlex, a novel conditional generative framework capable of generating multiple diverse CFs in a single forward pass. DiCoFlex is model-agnostic and does not require constant access to the underlying predictive model at any stage, relying solely on a dataset labeled by a classification model. Our approach leverages conditional normalizing flows to learn the distribution of counterfactual explanations that inherently satisfy user-specified constraints.
In contrast to existing methods, DiCoFlex incorporates customizable constraints such as sparsity and actionability directly at inference time, eliminating the need for retraining when constraints change. During inference, DiCoFlex achieves computational efficiency by generating multiple diverse CFs through a single forward pass, thus avoiding the computational burden inherent in methods that require separate optimization procedures for each counterfactual. As illustrated in Figure 1, our method can operate under different constraints. DiCoFlex generates diverse and plausible counterfactual explanations without any restrictions imposed on sparsity or actionability (as shown in Figure 1(a)). Sparsity, controlled by the user-defined exponent $p$ in $L_{p}$ norm, ensures minimal modifications in input features (as demonstrated in Figure 1(b)), while actionability constraints allow users to specify immutable features through a mask, guiding the generation towards practically feasible explanations (shown in Figure 1(c)). These parameters can be dynamically adjusted at inference time, providing unprecedented flexibility in generating counterfactual explanations that satisfy domain-specific constraints while maintaining a balance between diversity and validity without requiring explicit diversity penalties.
The main contributions of our work are summarized as follows:
  • We propose DiCoFlex, a classifier-agnostic, conditional generative framework that generates multiple counterfactuals efficiently in a single forward pass, without the need for model retraining or optimization per query instance.
  • Our method enables real-time customization of constraints including sparsity and actionability, providing practical, user-centric control over counterfactual generation.
  • We extensively evaluate DiCoFlex on benchmark datasets, demonstrating superior performance in terms of validity, proximity, diversity, and constraint compliance compared to existing state-of-the-art approaches.
2 Related Work
Counterfactual explanations (CFs) have become integral to explainable AI by providing intuitive alternatives that elucidate model decisions [40]. Optimization-based methods formulate CF generation as minimizing distance between instances while ensuring prediction changes, using implementations including gradient-based approaches [40], integer programming [37], genetic algorithms [5], and satisfiability problems [15]. Instance-based methods like FACE [26] select examples yielding desired predictions from existing data, while CBCE [16] merges features from paired contrasting instances. Model-specific approaches such as decision tree methods [11, 35] leverage model structure by traversing alternative paths.
Diverse Counterfactual Explanations
Generating multiple diverse counterfactuals offers significant advantages over single explanations by providing alternative paths for recourse and better illustrating decision boundaries [17, 19]. DiCE [19] directly incorporates diversity into its loss function but requires separate optimization for each explanation, increasing computational costs. Other approaches achieve diversity through various strategies: DiVE [30] employs determinantal point processes, while methods like [36, 32, 4] partition the feature space. Multi-objective approaches like MOC [5] generate solutions representing different trade-offs between properties such as proximity and sparsity, while CARE [27] and OrdCE [14] incorporate user preferences into the generation process.
Generative Models for Counterfactuals
Generative models have gained popularity for counterfactual generation due to their ability to produce more plausible explanations that remain on the data manifold. VAE-based approaches include ReViSE [13], C-CHVAE [25], CLUE [1], and CRUDS [8], which learn latent representations of the data distribution. GAN-based methods such as GeCo [33] and SCGAN [42] use adversarial training to balance validity and realism. Recent transformer-based approaches like TABCF [22] specifically target tabular data with mixed types. However, many of these approaches still require access to model gradients, produce redundant explanations, or lack flexibility in handling user-defined constraints.
Handling Constraints in Counterfactual Generation
Practical counterfactual explanations must satisfy various constraints including sparsity (minimizing modified features) and actionability (ensuring feasible changes). CERTIFAI [34] employs genetic algorithms to enforce constraints, while MCCE [28] conditions on immutable features. Methods focused on actionable recourse, such as CSCF [20], account for downstream effects of feature modifications, while Russell [31] and Verma et al. [39] generate explanations representing different modification strategies. However, most existing approaches require retraining or reoptimization when constraints change, limiting their flexibility in real-world applications.
Limitations of Current Approaches
Current counterfactual methods face substantial limitations for tabular data: optimization-based approaches require high computational costs for diverse explanations [19], and generative models often produce redundant explanations, require model gradients, or lack constraint flexibility [25]. Our proposed approach, DiCoFlex, addresses these limitations through a conditional generative framework that efficiently produces diverse counterfactuals in a single forward pass, enables real-time constraint customization, operates without requiring model access, and naturally handles mixed feature types.
3 DiCoFlex Method
Given a classification model $h$ and an initial input example $\mathbf{x}_{0}$ within a $d$-dimensional real space $\mathbb{R}^{d}$, one seeks to determine a counterfactual instance $\mathbf{x}^{\prime}\in\mathbb{R}^{d}$. A counterfactual $\mathbf{x}^{\prime}$ represents a sample, which (i) is classified to desired class $y^{\prime}$ by the model $h$ (i.e. $h(\mathbf{x}^{\prime})=y^{\prime}$), (ii) lies close enough to $\mathbf{x}_{0}$ according to some distance measure, $d(\mathbf{x}_{0},\mathbf{x}^{\prime})$, and (iii) is plausible (in-distribution sample).
This problem formulation focuses on the deterministic extraction of counterfactuals, where, for a given example $\mathbf{x}_{0}$ and desired class $y^{\prime}$, the model returns a single counterfactual $\mathbf{x}^{\prime}$. In practice, generating more plausible candidates provides a wider spectrum of possible explanations satisfying validity, plausibility, and proximity constraints. Therefore, we postulate an approximate conditional distribution for counterfactuals for a given example $\mathbf{x}_{0}$ and the target class $p(\mathbf{x}^{\prime}|\mathbf{x}_{0},y^{\prime})$.
In this section, we first introduce a general objective function for a model generating counterfactual explanations. Next, we parametrize our model with normalizing flow and explain our strategy for scoring counterfactual candidates in a training phase. Finally, we introduce sparsity and actionability constraints and summarize the training algorithm.
3.1 The model objective
We consider the parameterized model $p_{\theta}(\mathbf{x}^{\prime}|\mathbf{x}_{0},y^{\prime})$ for the distribution approximation. Direct optimization of $p_{\theta}(\mathbf{x}^{\prime}|\mathbf{x}_{0},y^{\prime})$ is challenging, because we do not have access to the training ground-truth pairs composed of initial examples $\mathbf{x}_{0}$ and the corresponding counterfactuals $\mathbf{x}^{\prime}$ for class $y^{\prime}$. Therefore, we introduce conditional distribution $q(\mathbf{x}^{\prime}|\mathbf{x}_{0},y^{\prime})$ and we utilize the expected value of Kullback-Leibler divergence (KLD) between $p_{\theta}(\mathbf{x}^{\prime}|\mathbf{x}_{0},y^{\prime})$ and $q(\mathbf{x}^{\prime}|\mathbf{x}_{0},y^{\prime})$ as training objective:
$\mathcal{L}$ $=\mathbb{E}_{\mathbf{x}\sim p_{\text{data}},y^{\prime}\sim\boldsymbol{\pi}\setminus\{h(\mathbf{x})\}}\left[D_{\text{KL}}(q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})\parallel p_{\theta}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime}))\right]$ (1)
$=\mathbb{E}_{\mathbf{x},y^{\prime}}\mathbb{E}_{\mathbf{x}^{\prime}\sim q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})}\left[\log\frac{q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})}{p_{\theta}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})}\right]$ (2)
$=\mathbb{E}_{\mathbf{x},y^{\prime}}\mathbb{E}_{\mathbf{x}^{\prime}\sim q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})}\left[\log q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})-\log{p_{\theta}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})}\right],$ (3)
where $\mathbf{x}$ is an example from the data distribution $p_{\text{data}}$ (training example) and $y^{\prime}$ is sampled from categorical distribution $\boldsymbol{\pi}$ representing the prior over classes, excluding the current class of $\mathbf{x}$, predicted by $h(\cdot)$. The component $q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})$ does not depend on parameters $\theta$, therefore the final objective is:
$\mathcal{Q}=-\mathbb{E}_{\mathbf{x},y^{\prime}}\mathbb{E}_{\mathbf{x}^{\prime}\sim q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})}\left[\log{p_{\theta}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})}\right].$ (4)
We further elaborate on how $p_{\theta}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})$ is modeled and how sample from $q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})$.
3.2 Modeling predictive conditional distribution
Our approach aims at directly optimizing the conditional negative log-likelihood given by (4) with the data sampled from the distribution $q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})$. As a consequence, it requires access to the parametrized density function. Several techniques, such as Kernel Density Estimation (KDE) and Gaussian Mixture Models (GMM), can be used to model the conditional density function $p_{\theta}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})$. In this work, we propose using a conditional normalizing flow model [29] for this purpose. Unlike KDE or GMM, normalizing flows do not rely on a predefined parametric form of the density function and are well-suited for modeling high-dimensional data. Furthermore, in contrast to other generative models, such as diffusion models or GANs, normalizing flows allow for exact density evaluation via the change of variables formula and can be efficiently trained by minimizing the negative log-likelihood. The general formula that represents normalizing flows can be expressed as:
$p_{\theta}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})=p_{Z}(f_{\theta}(\mathbf{x}^{\prime};\mathbf{x},y^{\prime}))\cdot\left|\det J_{f_{\theta}}(\mathbf{x}^{\prime};\mathbf{x},y^{\prime})\right|,$ (5)
where $f_{\theta}(\mathbf{x}^{\prime};\mathbf{x},y^{\prime})$ is an invertible transformation conditioned on $\mathbf{x}$ and $y^{\prime}$, $p_{Z}(f_{\theta}(\mathbf{x}^{\prime};\mathbf{x},y^{\prime}))$ is the base distribution, typically Gaussian and $J_{f_{\theta}}(\mathbf{x}^{\prime};\mathbf{x},y^{\prime})$ is Jacobian of the transformation. The biggest challenge in normalizing flows is the choice of the invertible function for which the determinant of Jacobian is easy to calculate. Several solutions have been proposed in the literature to address this issue with notable approaches, including NICE [6], RealNVP [7], and MAF [23]. Based on superior performance on benchmark dataset, we selected MAF as our normalizing flow model (see Appendix C).
3.3 Scoring training counterfactual explanations
The distribution $q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})$ is essential to sample data to train the flow model. Since we do not have direct access to this distribution, we introduce a score function, which quantifies what "the adequate counterfactual" means. Following the basic postulates defined by [10], the counterfactual has to satisfy at least validity, proximity, and plausibility constraints. These requirements lead us to the following score function:
$s(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})=\log p(y^{\prime}|\mathbf{x}^{\prime})-\lambda_{1}d(\mathbf{x},\mathbf{x}^{\prime})+\lambda_{2}p_{\text{data}}(\mathbf{x}^{\prime}),$ (6)
where $p(y^{\prime}|\mathbf{x}^{\prime})$ is class probability returned by model $h(\cdot)$, $p_{\text{data}}(\mathbf{x}^{\prime})$ is value of density function for $\mathbf{x}^{\prime}$ and $\lambda_{1},\lambda_{2}$ are user-defined trade-off parameters. Making use of the above criterion, we can evaluate the quality of a counterfactual candidate.
There are several approaches for including the above score function to the learning process of normalizing flow modeling the distribution of counterfactual explanations. However, to keep our model as simple as possible, we resort to the straightforward technique of querying training examples, which are similar to the input $\mathbf{x}$ and satisfy $h(\mathbf{x}^{\prime})=y^{\prime}$. For this purpose, we use a typical KNN (K-nearest neighbors) heuristic, in which we sample $K$ data points labeled as $y^{\prime}$ by the classifier that are closest to the input $\mathbf{x}$. Since we are sampling training examples from class $y^{\prime}$, the validity and plausibility constraints hold, while the proximity is satisfied by using the KNN approach.
Concluding, the approximation for the probability distribution $q(\cdot)$ is as follows:
[…]
where $N(\mathbf{x},y^{\prime},K,d)$ is set of $K$ closest neighbors of $\mathbf{x}$ in class $y^{\prime}$, considering distance measure $d(\cdot,\cdot)$.
One could ask whether a direct use of the distribution $\hat{q}(\mathbf{x}^{\prime}\mid\mathbf{x},y^{\prime},d)$ can be an alternative for generating counterfactual explanations, instead of learning a parametrized flow-based model ${p_{\theta}(\mathbf{x}^{\prime}\mid\mathbf{x},y^{\prime})}$. Observe that the use of $\hat{q}$ is inherently limited by the number of available training examples and lacks the ability to generalize to unseen instances. Additionally, the diversity of generated examples is constrained by the number of selected neighbors. Modeling a generative model provides smooth changes over input examples and allows for extrapolating counterfactual generation outside training data.
3.4 Sparsity on continuous features
In many practical use cases, counterfactuals need to satisfy additional constraints, such as sparsity. This means that, in addition to minimizing the Euclidean distance, a counterfactual has to minimize the number of modified attributes. While sparsity is often implicitly achieved for categorical variables due to the applied optimization scheme, enforcing sparsity on continuous features remains challenging. We show that the sparsification of continuous attributes can be controlled in the inference phase by an auxiliary conditioning factor in a flow model.
Instead of querying neighbor examples to $x$ according to Euclidean distance in KNN search, we use $L_{p}$ distance function given by:
$d_{p}(\mathbf{x}_{i}^{\prime},\mathbf{x})=||\mathbf{x}_{i}^{\prime}-\mathbf{x}||_{p}=\left(\sum_{j=1}^{D}|x^{\prime}_{j}-x_{j}|^{p}\right)^{1/p}.$ (8)
We can observe that the level of modified attributes can be controlled through an appropriate choice of the value of $p$. For instance, if $p=0.01$, neighbors with fewer modified attributes are favored, thereby enforcing sparsity. Making use of $p=2$ gives a standard Euclidean proximity measure. Since categorical attributes are commonly encoded by one-hot vectors, the above scheme affects mostly continuous features.
In practice, the sparsity level can be effectively controlled in the inference time by incorporating the parameter $p$ into the flow model $p_{\theta}(\mathbf{x}^{\prime}\mid\mathbf{x},y^{\prime},p)$ through additional conditioning. During training, values of $p$ are sampled and explicitly included in the conditioning of the model, allowing it to learn the relationship between $p$ and the desired sparsity in the generated outputs. As a result, during the inference stage, users can adjust the sparsity level of the modified attributes simply by selecting an appropriate value of $p$. This mechanism provides a flexible and interpretable way to regulate the extent of changes applied to the input $\mathbf{x}$, depending on specific needs or constraints. Note, however, that the applied $L_{p}$ norm does not freeze the continuous attributes but only encourages small changes on their portion, which can be seen as a type of soft constraint.
3.5 Actionability constraints
One of the crucial aspects of counterfactual explanations is enforcing the manipulation of some particular attributes. In order to accomplish that, in our approach, we introduce a further modification in a distance function while calculating the nearest neighbors:
$d_{p,\mathbf{m}}(\mathbf{x},\mathbf{x}^{\prime})=\|\mathbf{x}-\mathbf{x}^{\prime}\|_{p,m}=\alpha\sum_{j=1}^{D}m_{j}|x_{j}-x^{\prime}_{j}|^{p}+\sum_{j=1}^{D}(1-m_{j})|x_{j}-x^{\prime}_{j}|^{p},$ (9)
where $\mathbf{m}\in\{0,1\}^{D}$ is binary vector, for which $m_{d}=1$ indicates the attributes that should not be modified. The value of hyperparameter $\alpha$ should be large enough to prevent attributes $m_{d}=1$ from modification.
During training the masking vectors $\mathbf{m}$ are sampled from some set of possible mask patterns $\mathcal{M}$ and the mask is also delivered as a conditioning factor to the flow-based model, which represents $p_{\theta}(\mathbf{x}^{\prime}\mid\mathbf{x},y^{\prime},p,\mathbf{m})$. Once trained with a set of masks $\mathcal{M}$, DiCoFlex can dynamically change the mask during inference resulting in imposing different user expectations. Similar to enforcing sparsity, here we do not prevent from modifying masked attributes completely but only encourage the model to prioritize changing unmasked features first.
3.6 Training
Concluding, the training procedure for estimating $\theta$ is as follows. For each training iteration, $\mathbf{x}$ is selected from data examples and the target class $y^{\prime}$ is sampled for counterfactual explanation. Value $p$ that represents the sparsity level is also sampled from some predefined set $\mathcal{P}$, as well as the masking vector $\mathbf{m}$ from $\mathcal{M}$. Next, the set $N(\mathbf{x},y^{\prime},d_{p,\mathbf{m}},K)$ of $K$ nearest neighbors is calculated considering distance measure $d_{p,\mathbf{m}}$ given by (9). The counterfactual $\mathbf{x}^{\prime}$ is sampled from $\hat{q}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime},d_{p,\mathbf{m}})$ given by (7). Finally, $\theta$ is updated in gradient-based procedure by optimizing (4), considering conditioning both on $p$ and m. Appendix D contains the details of the training algorithm.
4 Experiments
In this section, we conduct a series of experiments that evaluate the quality of counterfactual explanations generated by various methods. Our main objective is to understand how effectively these counterfactuals flip the predictions of a model while maintaining realism, requiring minimal changes to the input, and offering diversity among the generated alternatives.
4.1 Experimental Setup
Datasets
We evaluate all methods on five well-established benchmark datasets commonly used in counterfactual explanation research: Adult [2] (income prediction), Bank Marketing [18] (customer response classification), Default [41] (credit card default risk prediction), Give Me Some Credit (GMC) [25] (financial insolvency prediction), and Lending Club [12](loan creditworthiness classification), see Appendix A for details. These datasets are selected for their diversity in dimensionality, feature types (mixed categorical and continuous variables), and real-world applicability, allowing us to evaluate the robustness and generalizability of our approach across different domains.
Parameterization of DiCoFlex
For our experimental evaluation, we compare two variants of our method: DiCoFlex with default Euclidean distance and s-DiCoFlex which uses $L_{p}$ norm with $p=0.01$. This latter version encourages lower sparsity for continuous features, leading to explanations that alter fewer continuous features while maintaining validity, see Section 3.5.
Baseline Methods
We compare our approach against several state-of-the-art counterfactual explanation methods. First of all, we consider methods, which generate multiple explanations for a single data point. CCHVAE [25] employs a conditional variational autoencoder to generate counterfactuals consistent with the data distribution. DiCE [19] uses gradient-based optimization to produce diverse counterfactuals in a model-agnostic framework. For all these methods, we generate up to 100 counterfactual explanations for each factual instance whenever feasible and randomly select 10 samples based solely on the validity criterion to provide an equal number of counterfactuals for every method.
Additionally, we compare with methods focused on returning a single explanation. ReViSE [13] introduces a probabilistic framework designed to separate sensitive and non-sensitive features with a focus on interpretability. The approach by Wachter et al. [40] represents a pioneering optimization-based method that minimizes a loss function balancing proximity to the input with the goal of changing the prediction. TABCF [22] specifically targets tabular data with a transformer-based VAE approach. This class of methods is expected to perform better because they are optimized for producing a single best explanation while the first class of methods takes a diversity criterion as a priority.
Evaluation Metrics
We evaluate counterfactual quality using metrics aligned with key desirable properties. For test instances originally classified as class 0, we generate corresponding counterfactuals targeting class 1 and vice versa. Validity measures the percentage of counterfactuals that successfully change model predictions to the target class. Hypervolume [21] is used to measure diversity to quantify the spread and coverage of counterfactuals in the objective space. Plausibility is assessed through Local Outlier Factor (LOF) [3] that measure isolation from the training distribution. Sparsity quantifies feature modifications, calculated as the proportion of changed features in each category. For continuous features, we use $\varepsilon$-sparsity, which counts features modified beyond a threshold of 5% of the feature range. Proximity measures distance between original instances and the counterfactuals using the Euclidean distance only for continuous features. Observe that for categorical features proximity coincides with the sparsity. Probability measures model confidence in the target prediction. For the sparsity, proximity and plausibility metrics, lower values indicate better performance ($\downarrow$), while for the validity, probability and diversity metrics, higher values are preferable ($\uparrow$). We provide a detailed description of the metrics in Appendix B.
4.2 Evaluating multiple explanations
Table 1: Comparison of counterfactual generation methods across datasets. The best method, which generates multiple counterfactual candidates for each metric is highlighted in bold (DiCoFlex, CCHVAE, DiCE). If there is a method that generates a single counterfactual (ReViSE, TABCF, Wachter) that is better than the bold result, it is underlined.
Classif. Proximity Sparsity $\epsilon$-sparsity LOF Hypervol.
Dataset Model Validity. ↑ prob. ↑ cont. ↓ cat. ↓ cont. ↓ log scale ↓ log scale ↑
Lending DiCoFlex 1.000 0.999 0.668 0.412 0.840 0.476 13.118
Club s-DiCoFlex 1.000 0.997 0.704 0.635 0.737 0.453 11.373
CCHVAE 1.000 0.682 0.769 0.137 0.895 0.519 7.448
DiCE 1.000 0.978 3.444 0.146 0.840 0.567 8.252
ReViSE 0.910 0.721 0.588 0.077 0.639 0.040 -
TABCF 1.000 0.946 0.557 0.635 0.551 0.047 -
Wachter 0.950 0.530 1.268 0.000 0.696 0.943 -
Adult DiCoFlex 1.000 0.998 0.581 0.515 0.498 2.156 2.036
s-DiCoFlex 1.000 1.000 0.373 0.568 0.350 1.812 0.379
CCHVAE 1.000 0.894 0.616 0.302 0.582 2.176 1.571
DiCE 1.000 1.000 5.096 0.557 0.582 3.815 0.681
ReViSE 0.860 0.879 1.004 0.206 0.576 2.637 -
TABCF 1.000 0.986 3.129 0.263 0.348 3.746 -
Wachter 0.970 0.578 1.046 0.000 0.830 2.893 -
Bank DiCoFlex 1.000 0.971 0.717 0.499 0.786 0.276 3.808
s-DiCoFlex 1.000 0.894 0.584 0.480 0.673 0.247 3.006
CCHVAE 1.000 0.951 0.989 0.357 0.816 0.778 0.309
DiCE 1.000 0.994 7.117 0.106 0.787 0.839 0.323
ReViSE 0.230 0.823 1.119 0.285 0.615 0.320 -
TABCF 1.000 0.850 3.747 0.210 0.441 0.360 -
Wachter 0.830 0.530 3.826 0.000 0.962 0.887 -
Default DiCoFlex 1.000 0.986 0.459 0.419 0.831 0.038 62.427
s-DiCoFlex 1.000 0.988 0.489 0.514 0.787 0.034 61.593
CCHVAE 1.000 0.963 0.679 0.463 0.832 0.047 53.706
DiCE 1.000 0.914 1.039 0.263 0.692 0.373 51.184
ReViSE 0.170 0.784 0.595 0.281 0.613 0.033 -
TABCF 1.000 0.884 0.516 0.510 0.412 0.057 -
Wachter 0.930 0.557 4.494 0.000 0.976 0.658 -
GMC DiCoFlex 1.000 0.969 0.690 0.755 0.765 0.861 15.667
s-DiCoFlex 1.000 0.968 0.781 0.764 0.763 0.899 15.810
CCHVAE 1.000 0.932 0.389 0.756 0.762 0.564 9.761
DiCE 1.000 0.934 3.478 0.753 0.699 0.399 9.184
ReViSE 0.780 0.909 8.525 0.180 0.877 0.702 -
TABCF 0.130 0.630 0.521 0.359 0.835 0.303 -
Wachter 0.950 0.584 8.588 0.000 0.987 2.436 -
The comparison with baselines for diverse counterfactual explanations (DiCE, CCHVAE) presented in Table 1 confirms that DiCoFlex returns more diverse explanations than competitive methods, achieving superior hypervolume scores. In terms of plausibility evaluation, both versions of DiCoFlex consistently outperform other baseline methods as indicated by the LOF measure (except the GMC dataset). Furthermore, the results reveal that the proposed method is capable of producing counterfactual explanations with a low proximity score on continuous attributes.
However, impressive diversity, plausibility, and proximity come at the price of slightly worse sparsity. Although s-DiCoFlex obtained the lowest $\epsilon$-sparsity on continuous features in three out of five cases, the categorical sparsity is higher than the one returned by DiCE and CCHAVE. It is apparent that these two methods switch categorical variables less often than continuous ones, which is usually caused by the applied optimization scheme and can be seen as their drawback. Observe that the use of s-DiCoFlex allows us to balance the level of sparsity between categorical and continuous features.
Our analysis confirms that typical measures for evaluating counterfactual explanations are in conflict and that there does not exist a method that optimizes all these criteria simultaneously. However, the proposed method compares favorably with other approaches beating them in diversity, plausibility, and proximity on continuous features.
4.3 DiCoFlex vs. s-DiCoFlex
As mentioned, lowering the value of $p$ in the $L_{p}$ norm for DiCoFlex directly results in lowering the sparsity measure for continuous attributes. As a side effect, it allows for larger changes in values of categorical variables to produce valid counterfactuals. An interesting consequence is that s-DiCoFlex generates more plausible samples than DiCoFlex (see the LOF score), which may be attributed to the way we model the distribution $q$ used for scoring counterfactual candidates; see Section 3.3. For low $p=0.01$, we restrict the support of $q$ to a small region of data space, where only a small number of coordinates are likely to be modified. Increasing this value leads to neighbors located in a larger and more diverse region since we allow for modifying all coordinates without extra penalty. Consequently, DiCoFlex with $p=2$ can sample out-of-distribution counterfactuals more often than s-DiCoFlex with $p=0.01$. In practice, dynamic change of $p$ in inference allows users to examine multiple counterfactual candidates and select the one that suits their expectations best.
4.4 Comparison with single counterfactual methods
To present a broader view on the performance of DiCoFlex, we include an additional comparison with commonly used methods that generate a single explanation (ReViSE, TABCF, Wachter) (see Table 1). We note that DiCoFlex and s-DiCoFlex still demonstrate impressive performance in terms of LOF score in this comparison. They both achieve the lowest LOF on Adult and Bank datasets and are comparable to the best-performing ReViSE method on the Default dataset. Moreover, they obtained the best continuous proximity measure in 3 out of 5 cases. This means that even though DiCoFlex produces multiple diverse counterfactuals for each instance, most of them satisfy the highest requirements used in counterfactual analysis.
Among competitive methods, ReViSE and TabCF represent strong baselines that achieve a satisfactory level of plausibility and sparsity. Nevertheless, these models often fail to generate valid counterfactuals (see TABCF on GMC or ReViSE on Bank and Default), which partially limits their usefulness. Generating multiple diverse counterfactuals (as in our method) allows the user to select suitable candidates that meet the expected criteria. It is interesting that TABCF and ReViSE obtain more than 10 times lower LOF on the Lending Club dataset than other approaches. It might follow from the fact that it is impossible to generate multiple good counterfactuals for this dataset, and only a few nearby candidates represent in-distribution samples. We also note that the classical Wachter method does not change categorical variables at all and generally gives suboptimal results, except for the validity score.
4.5 Runtime of counterfactuals generation
Figure 2: Average runtime (log scale).
A key advantage of our approach is that once trained, DiCoFlex can be used to generate multiple counterfactual explanations in a single forward pass. To illustrate this gain, we measure the time every method needs to generate a single counterfactual for 100 data points. Figure 2 confirms that both variants of DiCoFlex achieve generation times of $0.12\pm 0.05$ seconds, outperforming other baseline methods by orders of magnitude. Optimization-based methods demonstrate substantially higher computational costs, with DiCE, ReViSE, and TABCF all requiring over 1000 seconds on average. This efficiency enables real-time exploration of diverse counterfactual explanations with dynamic constraint adjustment, addressing a critical limitation of existing approaches.
4.6 Actionability Control
DiCoFlex is capable of imposing additional constraints when generating counterfactuals without retraining the model. As demonstrated in Section 4.3 and Table 1, the use of the $L_{p}$ norm with small $p$ allows to minimize the modifications of the continuous features and improve plausibility. In this experiment, we show the effect of introducing actionability constraints, which prevents severe modification of user-defined features.
Table 2: Comparison of the unconstrained version of DiCoFlex with its masked variant, evaluated on features selected by the given mask. We report $\epsilon$-sparsity for numerical variables and standard sparsity for categorical ones (lower score means less attributes affected).
mask 1 mask 2 mask 3 mask 4
Method Capital Gain Capital Loss Age Race Sex Native Country
Masked-DiCoFlex 0.023 0.037 0.069 0.052 0.045 0.030
Standard DiCoFlex 0.148 0.097 0.833 0.170 0.200 0.116
For illustration, we train DiCoFlex on the Adult dataset with four mask configurations, which cover one or two features, as shown in Table 2. For instance, the use of "mask 1" should block changes of "Capital Gain" and "Capital Loss" attributes. As can be observed in the result, the application of every individual mask indeed decreases the sparsity level compared to the unconstrained version when no mask is used. In consequence, this flexible control enables users to dynamically adjust both sparsity and actionability, generating counterfactuals satisfying different domain-specific requirements without requiring model retraining. Appendix G contains additional evaluation of these models.
5 Conclusion
We presented DiCoFlex, a model-agnostic conditional generative framework for counterfactual explanations that addresses key limitations in existing approaches. By leveraging normalizing flows, our method enables counterfactual generation without constant access to the classification model, provides inference-time control over constraints, and efficiently produces multiple diverse explanations in a single forward pass. Empirical evaluations on five benchmark datasets demonstrate that DiCoFlex outperforms the state-of-the-art methods in diversity and validity while achieving superior plausibility and computational efficiency. Although optimization-based methods may achieve better sparsity in some cases, our approach offers a more balanced trade-off between competing explanation criteria.
Appendix A Dataset Descriptions
This appendix provides detailed descriptions of the datasets used in our experimental evaluation. Table 3 summarizes the key characteristics of each dataset, including the number of samples used for our experiments, number of numerical and categorical features, and number of classes.
Table 3: Dataset characteristics and statistics.
Dataset Samples # Feat # Num # Cat Target Class Dist.
Lending Club 30,000 12 8 4 21.8% default
Give Me Some Credit 30,000 9 6 3 6.7% yes
Bank Marketing 30,000 16 7 9 11.3% yes
Credit Default 27,000 23 14 9 22.1% yes
Adult Census 32,000 12 4 8 24.5% $\geq$50K
A.1 Lending Club
The Lending Club dataset [12] contains detailed information about loans issued through the Lending Club peer-to-peer lending platform. It includes borrower characteristics (such as credit score, annual income, employment length), loan specifics (loan amount, interest rate, purpose), and performance indicators (payment status, delinquency). The binary classification task is to predict whether a loan will be fully paid or charged off (default). This dataset is particularly relevant for financial counterfactual explanations as it represents real-world credit risk assessment scenarios where understanding model decisions is crucial for both borrowers and lenders.
A.2 Give Me Some Credit
The Give Me Some Credit dataset [25] contains anonymized records of credit users with features such as debt-to-income ratio, number of times delinquent, monthly income, age, and number of open credit lines. The target variable indicates whether a user experienced a serious delinquency (more than 90 days overdue) within the previous two years. This dataset provides insight into credit risk prediction in consumer finance, where counterfactual explanations can offer actionable guidance to consumers looking to improve their creditworthiness.
A.3 Bank Marketing
The Bank Marketing dataset [18] contains information from a direct marketing campaign conducted by a Portuguese banking institution. The features include client data (age, job, marital status, education), campaign contact information (communication type, day, month), economic indicators, and previous campaign outcomes. The prediction task is to determine whether a client will subscribe to a term deposit. This dataset represents a real-world marketing scenario where understanding model decisions can improve campaign efficiency and provide insights for personalized marketing strategies.
A.4 Credit Default
The Credit Default dataset [41] contains information on credit card clients in Taiwan, including demographic factors, credit data, payment history, and bill statements. The target variable indicates whether the client defaulted on their payment in the following month. With 23 features (14 numerical and 9 categorical), this dataset presents complex feature interdependencies common in financial data. The dataset is valuable for counterfactual explanation research because it represents real-world credit risk assessment with diverse feature types and nonlinear relationships.
A.5 Adult
The Adult Census dataset [2] contains demographic information extracted from the 1994 U.S. Census database. Features include age, education, occupation, work hours per week, and capital gain/loss. The binary classification task is to predict whether an individual’s income exceeds \$50,000 per year. This dataset is widely used in fairness and explainability research, as it contains sensitive attributes like race, gender, and age, making it valuable for studying how counterfactual explanations handle demographic factors.
Appendix B Evaluation Metrics
We evaluate counterfactual quality using metrics that correspond to key desiderata. For a test set $\mathcal{X}^{0}_{\text{test}}=\{\mathbf{x}^{0}_{n}|h(\mathbf{x}^{0}_{n})=0\}_{n=1}^{N}$, we generate set of $N$ counterfactuals $\mathcal{X^{\prime}}^{1}=\{\mathbf{x}^{\prime 1}_{n}\}_{n=1}^{N}$ for class $1$ with the evaluated model.
Validity measures the success rate of changing model predictions:
$\text{Validity}\ (\uparrow)=\frac{N_{\text{val}}}{N}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}(h(\mathbf{x}^{\prime}_{n})=1)$ (10)
Sparsity quantifies feature modifications, separately for categorical and numerical features:
[…]
where $\mathbf{x}^{0}_{n,\text{cat/num}}$ represents $n$-th generated counterfactual reduced to categorical numerical attributes.
For continuous features, we use $\epsilon$-sparsity, counting features modified beyond $\epsilon\cdot V$ ($V$ is feature range, $\epsilon=0.05$):
$\epsilon\text{-Sparsity cont.}\ (\downarrow)=\frac{1}{N_{\text{val}}}\sum_{n=1}^{N_{\text{val}}}\frac{1}{D}\sum_{d=1}^{D}\mathbb{I}(|x^{0}_{n,d}-x^{\prime 1}_{n,d}|>\epsilon\cdot V_{g})$ (12)
Proximity measures distance to original instances using Gower distance for mixed data and Euclidean for continuous features:
Proximity $\text{Proximity Num}\ (\downarrow)=\frac{1}{N_{\text{val}}}\sum_{n=1}^{N_{\text{val}}}\|\mathbf{x}^{0}_{n,\text{num}}-\mathbf{x}^{\prime 1}_{n,\text{num}}\|_{1}$ (13)
Predictive Performance is assessed via Validity and Classif. prob. (model confidence in target prediction).
Diversity is measured by Hypervol. log scale [21], which quantifies spread and coverage in objective space.
Plausibility is evaluated with LOF log scale (isolation from training distribution) and Log Density (probability under training data distribution).
Lower values are better for sparsity, proximity, and plausibility metrics ($\downarrow$), while higher values are better for validity, model confidence, and diversity ($\uparrow$).
Appendix C Generative Model Selection
We conducted an ablation study to select the most suitable normalizing flow architecture for our framework, comparing different models based on negative log-likelihood (NLL) on the Adult dataset.
Table 4: Negative log-likelihood comparison of generative models on the Adult dataset. Lower values indicate better performance.
MAF NICE RealNVP KDE
NLL -43.2998 26.6644 26.5827 30.5120
We compared three normalizing flow architectures—MAF [23], NICE [6], and Real NVP [7], as well as KDE as a non-parametric baseline. MAF significantly outperformed all alternatives, while NICE and RealNVP showed comparable performance, and KDE exhibited the poorest results.
MAF’s superior performance stems from its autoregressive structure, which enables more expressive transformations by conditioning each dimension on previously transformed dimensions. This property is crucial for accurately modeling the conditional distribution of counterfactual explanations. Based on these results, we selected MAF as the architecture for DiCoFlex, contributing significantly to its ability to generate high-quality, diverse counterfactual explanations.
Appendix D Details of training algorithm
Algorithm 1 Training procedure
Require: number of steps $T$, training examples $\mathcal{X}$, classification model $h(\cdot)$, prior class distribution $\boldsymbol{\pi}$, set of sparsity levels $\mathcal{P}$, set of considered masking $\mathcal{M}$, number of nearest neighbors $K$
Initialize $\theta_{0}$
for $t=1$ to $T$ do
     Sample $\mathbf{x}\sim\mathcal{X}$ and class $y^{\prime}\sim\boldsymbol{\pi}\setminus\{h(\mathbf{x})\}$
     Sample sparsity level $p\sim\mathcal{P}$ and mask $m\sim\mathcal{M}$;
     Sample $K$ counterfactuals $\mathbf{x}^{\prime}\sim\hat{q}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime},d_{p,\mathbf{m}})$ given by (7);
     Update parameters $\theta_{t}$ by optimizing $\mathcal{Q}$ given by (4);
end for
return $x_{0}$
The training procedure is outlined in Algorithm 1. In each training iteration, a data example $\mathbf{x}$ is drawn from the dataset $\mathcal{X}$, and a target class $y^{\prime}$ is sampled from the class distribution excluding the current class of $\mathbf{x}$, i.e., $\boldsymbol{\pi}\setminus{h(\mathbf{x})}$. Subsequently, a sparsity level $p$ is selected from a predefined set $\mathcal{P}$, along with a corresponding masking vector $\mathbf{m}$ from the set $\mathcal{M}$.
Next, the set of $K$ nearest neighbors, denoted by $N(\mathbf{x},y^{\prime},d_{p,\mathbf{m}},K)$, is computed using the distance measure $d_{p,\mathbf{m}}$ defined in (9). A counterfactual example $\mathbf{x}^{\prime}$ is then sampled from the distribution $\hat{q}(\mathbf{x}^{\prime}\mid\mathbf{x},y^{\prime},d_{p,\mathbf{m}})$ as defined in (7). Finally, the parameters $\theta_{t}$ are updated via a gradient-based optimization procedure aimed at maximizing the objective in (4), with conditioning on both $p$ and $\mathbf{m}$.
Appendix E Limitations of our method
DiCoFlex is deliberately designed for tabular data with mixed feature types, the most common domain for counterfactual explanations in high-stakes decision-making contexts. Although this focus enables strong performance where interpretability is most needed, extending to other data modalities would require architectural adaptations. Our approach exhibits natural trade-offs between competing explanation criteria, favoring diversity and plausibility while maintaining competitive performance in sparsity metrics. The computational efficiency during inference requires an initial training investment, though this one-time cost enables subsequent real-time applications that outperform existing methods by orders of magnitude.
Appendix F Computational Resources
Our experimental framework utilized Python [38] as the primary programming language, Additionally, the open-source machine learning library PyTorch [24] is used to implement DiCoFlex. All experiments were conducted on a GPU cluster equipped with a GeForce RTX 4090 graphics card (24 GB VRAM) and an AMD Ryzen Threadripper PRO 5975WX 32-core processor, with 256 GB of available RAM. These resources provide sufficient computational power and processing speed to meet the requirements of our algorithm.
Appendix G Additional experimental results
G.1 Runtime Comparison Across Datasets
Figure 3 displays runtime comparisons between DiCoFlex and baseline methods across all datasets, with time shown on a logarithmic scale. Both variants of DiCoFlex consistently achieve substantially faster execution times for counterfactual generation, while competing approaches demand processing times that are orders of magnitude longer. This substantial performance advantage stems from fundamental architectural differences. DiCE [19] performs separate optimization procedures for each explanation. CCHVAE [25] requires expensive latent space searching. Similarly, Wachter [40], ReViSE [13] and TABCF [22] rely on iterative or gradient-based optimization procedures that scale poorly with the number of counterfactuals generated. In contrast, DiCoFlex leverages conditional normalizing flows trained solely on labeled data to generate multiple diverse counterfactuals in a single forward pass. By eliminating iterative optimization procedures and model access at inference time, DiCoFlex enables real-time counterfactual generation.
G.2 Impact of Actionability Constraints
Table 5 presents the evaluation of counterfactual explanations generated by DiCoFlex with different actionability constraints imposed through feature masks. The analysis reveals complex relationships between masking constraints and evaluation metrics. In particular, the application of different masks results in varying impacts across metrics without following a consistent directional pattern.
Mask 1, which prevents modifications to Capital Gain and Capital Loss, achieves the lowest continuous proximity score but exhibits reduced diversity, as indicated by its hypervolume score. In contrast, mask 4 (restricting Sex and Native Country modifications) yields the highest hypervolume while maintaining moderate performance across other metrics. Mask 3 (constraining Age) demonstrates high classification probability but the highest proximity score, indicating a greater deviation from the original instances.
The observed results indicate that actionability constraints introduce complex trade-offs that do not follow simple patterns. The absence of consistent correlations between metrics under different masking configurations suggests that performance characteristics are highly dependent on the specific constraints applied rather than adhering to predictable trade-off relationships.
These findings further validate DiCoFlex flexibility in selecting various user-defined constraints without retraining, while demonstrating that the selection of appropriate constraints should be guided by domain-specific requirements rather than general optimization principles. The mechanism enables the practical customization of counterfactual explanations according to specific application needs, where different feature restrictions may be necessary due to legal, ethical, or practical considerations.
(a) Adult
(b) Bank
(c) Default
(d) GMC
(e) Lending-Club
Figure 3: Visualization of runtime of DiCoFlex method and other baseline methods.
Table 5: Influence of imposing actionability constraints on the metrics used to evaluate counterfactual explanations.
Classif. Proximity Sparsity $\epsilon$-sparsity LOF Hypervol.
Model prob. ↑ cont. ↓ cat. ↓ cont. ↓ log scale ↓ log scale ↑
mask 1 0.987 0.496 0.557 0.483 1.881 0.881
msak 2 0.980 0.553 0.555 0.517 2.364 1.703
mask 3 0.993 0.627 0.568 0.513 2.310 2.411
mask 4 0.992 0.517 0.590 0.520 2.327 2.640
unconstrained 0.998 0.581 0.515 0.498 2.156 2.036